Use the compound interest formula to determine the interest earned in the given period.
P=$4000 at 6.5% compounded annually for 8 years
Please explain, thanks!

Answer by rfer(16322)

FV=4000(1.065)^8
FV=4000(1.655)
FV=10956.00
-------
10956-4000=6956 interest